# Superfluid Band Theory for the Rod Phase in the Magnetized Inner Crust Matter: Entrainment, Spin-orbit, Spin-triplet Pairing

### v1

@@ Abstract
The inner crust of neutron stars hosts a rich variety of nuclear phenomena and provides a unique environment for exploring microscopic nuclear properties relevant to diverse astrophysical observations. Particularly magnetars, which possess extremely strong magnetic-fields, have attracted increasing attention in connection with nuclear spin dynamics and unconventional pairing correlations. This work is dedicated to develop a comprehensive theoretical framework to describe the structures and properties of two-dimensional (rod-phase) matter in the neutron star inner crust, incorporating band-structure effects, neutron spin-triplet pairing, and strong magnetic-fields on an equal footing. Nuclear interactions and correlations are treated within the framework of nuclear density functional theory combined with a local description of pairing correlations, namely the superfluid local density approximation. Within this framework, the definition of the anomalous density is generalized to include spin-triplet components, and a spin-triplet pairing energy density functional is consistently incorporated.
The main results of this study can be summarized as follows. In the first place, the magnetic-fields of the order of $10^{16}\,$G are found to substantially enhance the neutron effective mass by a factor of approximately $1.5$, indicating a significant modification of entrainment properties in strongly magnetized crustal matter. In the second place, while the overall behavior of pairing phase transitions is qualitatively similar to that observed in one-dimensional systems studied previously, the present two-dimensional calculations reveal a nontrivial role of the spin-orbit interaction in inducing spin-polarization under magnetic fields. In the third place, concerning spin-triplet superfluidity, the rank-0 component is shown to emerge as a consequence of magnetic-field-induced spin-polarization, irrespective of the presence of spin-triplet pairing interactions, whereas the rank-2 component appears only when the corresponding interaction channel is included.
These findings provide new microscopic insights into superfluid and magnetic phenomena in the inner crust of neutron stars and offer a theoretical basis for future studies of magnetar dynamics and related astrophysical processes.
@@ 1 INTRODUCTION
Neutron stars are compact objects which serve as natural laboratories for probing the nuclear and hadron properties through the verification with the astronomical observations Özel and Freire (2016); Lattimer (2021). The inner crust region in a neutron star is filled with the Coulomb lattice of nuclei coexisting with free neutron gas Ravenhall et al. (1983); Hashimoto et al. (1984); Chamel and Haensel (2008); Kobyakov and Pethick (2013); Caplan and Horowitz (2017), where the neutron superfluidity plays an indispensable role Lombardo and Schulze (2001); Dean and Hjorth-Jensen (2003); Sedrakian and Clark (2019); Chamel (2024). The superfluid properties in the neutron-star crust are believed to have dominant impacts on various astrophysical phenomena associated with neutron stars, such as pulsar glitches and neutron-star cooling processes Link et al. (1999); Yakovlev and Pethick (2004); Peng et al. (2006); Shternin et al. (2011); Zhou et al. (2022). However, their detailed behavior can be altered radically by effects of the crystalline structures of nuclei including nuclear pasta phases, making a comprehensive analysis extremely challenging. For example, band-theoretical calculations applied to the inner-crust matter have suggested that the effective masses of free neutrons may be enhanced by factors of several up to around ten due to crystalline-structure effects Carter et al. (2005); Chamel (2005, 2012, 2017). This phenomenon, known as the entrainment effect, continues to be a subject of intense controversy, which involves a consistency with pulsar glitch interpretations Andersson et al. (2012); Chamel (2013); Haskell and Melatos (2015). Subsequently, calculations incorporating self-consistency and superfluidity have been performed Kashiwaba and Nakatsukasa (2019); Watanabe and Pethick (2017); Minami and Watanabe (2022), and a variety of methodologies have been proposed, including approaches in which the effective mass is extracted from the dynamical response through time evolutions Sekizawa et al. (2022); Yoshimura and Sekizawa (2024), as well as methods to determine the superfluid fraction from its static properties Almirante and Urban (2024b, a, 2025); Almirante et al. (2025). Within these researches regarding the neutron effective mass or superfluid fraction, it was often the case that only the $^{1}S_{0}$ component was taken into considerations. However, in the context of astronomical simulations, the impact of spin-triplet superfluid components, such as the $^{3}P_{2}$ component, has been actively investigated Chau et al. (1992); Schwenk and Friman (2004); Page et al. (2004, 2009, 2011); Shternin et al. (2011); Haskell and Melatos (2015); Potekhin et al. (2015); Masuda and Nitta (2016, 2020); Leinson (2020); Masaki et al. (2020); Andersson (2021); Kobayashi and Nitta (2023); Marmorini et al. (2024); Granados et al. (2025); Nam and Sekizawa (2025); Hattori and Sekizawa (2025); Sedrakian and Rau (2025), and these studies are pushing their physical interpretation to the next stage. Although spin-triplet components are generally expected to emerge only in the neutron-star core Sedrakian and Clark (2019); Wei et al. (2020), recent studies have begun to suggest their existence in the crustal matter or even in finite nuclei. For example, in Ref. Hinohara et al. (2024), it has been pointed out that, owing to the contribution of the spin-orbit interaction, $^{3}P_{1}$ superfluid components emerge in several open-shell nuclei. Another example is Ref. Tajima et al. (2023) where they revealed that spin-polarized neutron matter may produce the $^{3}P_{0}$ pairing component. There is a kind of neutron stars known as magnetars Makishima et al. (2014); Turolla et al. (2015); Kaspi and Beloborodov (2017), which possess surface magnetic fields exceeding $10^{14}\,$G and possibly reaching up to $10^{17}$ or $10^{18}$ G Potekhin and Yakovlev (1996); Broderick et al. (2000); Bonanno et al. (2003); Naso et al. (2008); Frieben and Rezzolla (2012). In such matter with enormous magnetic-fields, the nuclear compositions and superfluid phase structures may be drastically converted Muzikar et al. (1980); Broderick et al. (2000); Peña Arteaga et al. (2011); Chamel et al. (2012); Basilico et al. (2015); Stein et al. (2016a, b); Yasui et al. (2019); Parmar et al. (2023); Sekizawa and Kaba (2023); Jiang and Chen (2024); Basilico et al. (2025), which should be thoroughly investigated based on the microscopic theories.
@@
In our previous studies, under the title of “superfluid band theory,” we have pursued an approach in which superfluidity and band-structure effects are incorporated on the same footing into the conventional density functional theory (DFT) Hohenberg and Kohn (1964); Kohn and Sham (1965); Nakatsukasa et al. (2016); Colò (2020) for the superfluid systems Yoshimura and Sekizawa (2024). There has been a wealth of literature where the DFT calculations successfully elucidated structures and properties of the inner crust matter, including pasta phases Magierski and Heenen (2002); Gögelein et al. (2008); Newton and Stone (2009); Pais and Stone (2012); Grill et al. (2014); Schuetrumpf and Nazarewicz (2015); Fattoyev et al. (2017); Schuetrumpf et al. (2019); Pęcak et al. (2024). For the superfluid extension of the nuclear DFT, we employ a local treatment of pairing interactions, known as superfluid local density approximation (SLDA) Bulgac and Yu (2002); Bulgac (2002); Jin et al. (2021). By extending this framework into systems at finite temperatures and under finite magnetic-fields, we have further advanced the development of a tool to comprehensively describe the structures and properties of neutron star matter under extreme conditions, such as supernova matter and magnetars Yoshimura and Sekizawa (2025). In the present study, we extend our superfluid band theory to two-dimensional nuclear matter in the rod phase, which serves as a representative geometry of nuclear pasta. Our aim is to clarify how low-dimensional crystalline structures, in combination with spin-orbit interactions and strong magnetic-fields, influence neutron superfluidity, spin polarization, and pairing phases. To disentangle the roles of band-structure effects, magnetic fields, and pairing correlations, we adopt a stepwise approach: we firstly analyze the entrainment effect within the Hartree-Fock framework, and subsequently incorporate superfluidity to investigate the interplay between spin-polarization and spin-singlet pairing, under the effect of the spin-orbit interaction. We then extend the superfluid density functional framework to include spin-triplet pairing interactions, and systematically examine the emergence of spin-triplet superfluid components by performing calculations with and without these interactions under various magnetic-field strengths. Through this unified and self-consistent framework, this work aims to provide microscopic insights into superfluid and magnetic phenomena in neutron-star crust matter under extreme conditions.
@@
This article is arranged as follows. In the Section 2, the theoretical framework of this study is briefly explained, including the spin-triplet pairing. The Section 3 indicates the computational settings which are employed in this work. The calculation results are shown and discussed in Section 4. Section 5 summarizes the work and suggests the future prospects.
@@ 2 FORMALISM
In this section, we present the theoretical framework employed in the present study, starting from the generic formulations of the Skyrme Hartree-Fock (HF) and Hartree-Fock-Bogoliubov (HFB) theories. The formulation of band-structure effects within the superfluid framework has been developed in our previous work Yoshimura and Sekizawa (2024). For the spin-triplet pairing terms, detailed derivations can be found in Refs. Dobaczewski et al. (1984); Duguet (2014); Hinohara et al. (2024), although the notation used therein differs slightly from the present formulation. The extension to finite magnetic fields is based on the general formalisms discussed in Refs. Chamel et al. (2012); Stein et al. (2016a), and their implementation within our framework has been established in Ref. Yoshimura and Sekizawa (2025).
2.1 Skyrme-type Energy Density Functional
The total energy per nucleon in the neutron star matter is in general written by
@@ 2.1 Skyrme-type Energy Density Functional
where the nuclear part of the energy density is decomposed into kinetic, interaction, Coulomb, and (for superfluid systems) pairing contributions,
$\mathcal{E}_{\text{nucl}}(\bm{r})=\mathcal{E}_{\text{kin}}(\bm{r})+\mathcal{E}_{\text{int}}(\bm{r})+\mathcal{E}_{\text{Coul}}(\bm{r})\,(+\mathcal{E}_{\text{pair}}(\bm{r})).$ (2)
@@
$\mathcal{E}_{\text{kin}}(\bm{r})$ $=\sum_{q=n,p}\frac{\hbar^{2}}{2m_{q}}\tau_{q}(\bm{r})$ (3)
@@
where the exchange term of the Coulomb energy is evaluated using the Slater approximation. For the interaction part we adopt the Skyrme-type EDF which possesses the time-even and time-odd components, provided as
$\mathcal{E}_{\text{int}}(\bm{r})$ $=\mathcal{E}_{\text{even}}(\bm{r})+\mathcal{E}_{\text{odd}}(\bm{r})$ (5)
$\mathcal{E}_{\text{even}}(\bm{r})$ $=\sum_{t=0,1}\big[C^{\rho}_{t}\rho_{t}^{2}+C^{\alpha}_{t}\rho_{t}^{2}\rho_{0}^{\alpha}+C^{\Delta\rho}_{t}\rho_{t}\Delta\rho_{t}+C^{\tau}_{t}\rho_{t}\tau_{t}+C^{\nabla J}\rho_{t}\gradient\cdot\bm{J}_{t}-C^{T}_{t}\sum_{\mu,\nu=x}^{z}J_{t,\mu\nu}J_{t,\mu\nu}\big],$
$\mathcal{E}_{\text{odd}}(\bm{r})$ $=\sum_{t=0,1}\big[C^{s}_{t}\bm{s}_{t}^{2}+C^{s\alpha}_{t}\rho^{\alpha}_{0}\bm{s}_{t}^{2}+C^{T}_{t}\bm{s}_{t}\cdot\bm{T}_{t}+C^{\Delta s}_{t}\bm{s}_{t}\cdot\Delta\bm{s}_{t}-C^{\tau}_{t}\bm{j}_{t}\cdot\bm{j}_{t}+C^{\nabla J}\bm{s}_{t}\cdot(\gradient\times\bm{j}_{t}).$
The coupling constants $C_{t}^{X}$ are expressed in terms of the Skyrme force parameters, as summarized in Ref. Lesinski et al. (2007). Since the $\bm{s}_{t}\cdot\Delta\bm{s}_{t}$ term is known to induce spin instabilities Hellemans et al. (2012); Sekizawa and Yabana (2013), we set $C_{t}^{\Delta s}=0$ throughout this work. The local densities $\rho$, $\tau$, $\bm{j}$, $\bm{s}$, $\bm{T}$, and $J_{\mu\nu}$ denote the particle density, kinetic density, current density, spin density, spin-kinetic density, and spin-current tensor density, respectively. They are constructed from the single-particle wave functions as
@@
$J_{\mu\nu}(\bm{r})$ $=$ $\sum_{k,s,s^{\prime}}\frac{1}{2i}\quantity[\sigma_{\nu}]_{ss^{\prime}}\big[\psi^{*}_{ks^{\prime}}(\bm{r})\quantity(\nabla_{\mu}\psi_{ks}(\bm{r}))-\psi_{ks}(\bm{r})\quantity(\nabla_{\mu}\psi_{ks^{\prime}}(\bm{r}))\big].$ (11)
The spin-current tensor can be decomposed into its scalar (trace), vector, and traceless symmetric components as
@@
where the mean-field potentials are obtained from functional derivatives,
@@
$\frac{\delta E}{\delta J_{\mu\nu}}$ $=\mathsf{B}_{\mu\nu}(\bm{r})$
These potentials are explicit functionals of the local densities, and the ground state is obtained self-consistently by iterating the Kohn-Sham equations.

### v2

# Superfluid Band Theory for the Rod Phase in the Magnetized Inner Crust Matter: Entrainment, Spin-orbit Coupling, and Spin-triplet Pairing

@@ Abstract
The inner crust of neutron stars hosts a rich variety of nuclear phenomena and provides a unique opportunity for exploring microscopic nuclear properties relevant to diverse astrophysical observations. In particular, magnetars, which possess extremely strong magnetic-field, have attracted increasing interest in connection with nuclear spin dynamics and unconventional pairing correlations. This work is dedicated to develop a comprehensive theoretical framework to describe structures and properties of two-dimensional crustal matter (i.e., rod phase) in the inner crust of neutron stars, incorporating band-structure effects, neutron spin-triplet pairing, and strong magnetic-fields on an equal footing. Nuclear interactions and correlations are treated within the framework of nuclear density functional theory combined with a local description of pairing correlations, known as the superfluid local density approximation. Within this framework, the definition of the anomalous density is generalized to include spin-triplet components, and a spin-triplet pairing energy density functional is consistently incorporated. The main results of this study can be summarized as follows. In the first place, magnetic-fields of the order of $10^{16}\,$G are found to substantially enhance the neutron effective mass by a factor of approximately $1.5$, indicating a significant modification of entrainment properties in strongly magnetized crustal matter. In the second place, while the overall behavior of pairing phase transitions is qualitatively similar to that observed in one-dimensional systems studied previously, the present two-dimensional calculations reveal a nontrivial role of the spin-orbit interaction in inducing spin-polarization under magnetic-fields. In the third place, concerning the spin-triplet superfluidity, the rank-0 component is shown to emerge as a consequence of magnetic-field-induced spin-polarization, irrespective of the presence of spin-triplet pairing interactions, whereas the rank-2 component appears only when the corresponding interaction channel is included. These findings provide new microscopic insights into superfluid and magnetic phenomena in the inner crust of neutron stars and offer a theoretical basis for future studies of magnetars and related astrophysical processes.
@@ 1 INTRODUCTION
Neutron stars are compact objects which serve as natural laboratories for probing the nuclear and hadronic properties through the verification with the astronomical observations Özel and Freire (2016); Lattimer (2021). The inner crust region in a neutron star is composed of the Coulomb lattice of nuclei coexisting with free neutron gas Ravenhall et al. (1983); Hashimoto et al. (1984); Chamel and Haensel (2008); Kobyakov and Pethick (2013); Caplan and Horowitz (2017), where neutron superfluidity plays an indispensable role Lombardo and Schulze (2001); Dean and Hjorth-Jensen (2003); Sedrakian and Clark (2019); Chamel (2024). Superfluid properties in the neutron-star crust are believed to have impacts on various astrophysical phenomena associated with neutron stars, such as pulsar glitches and neutron-star cooling Link et al. (1999); Yakovlev and Pethick (2004); Peng et al. (2006); Shternin et al. (2011); Zhou et al. (2022). However, their detailed behavior can be altered radically by effects of the crystalline structures of nuclei including nuclear pasta phases, making a comprehensive analysis extremely challenging. For example, band-theoretical calculations applied to the inner-crust matter have suggested that the effective mass of free neutrons may be enhanced by factors of several up to around ten due to crystalline-structure effects Carter et al. (2005); Chamel (2005, 2012, 2017). This phenomenon, referred to as the entrainment effect, continues to be a subject of intense controversy, which involves consistency with pulsar glitch interpretations Andersson et al. (2012); Chamel (2013); Haskell and Melatos (2015). Subsequently, calculations incorporating self-consistency and superfluidity have been performed Kashiwaba and Nakatsukasa (2019); Watanabe and Pethick (2017); Minami and Watanabe (2022), and a variety of methodologies have been proposed, including approaches in which the effective mass is extracted from the dynamical response through time evolutions Sekizawa et al. (2022); Yoshimura and Sekizawa (2024), as well as methods to determine the superfluid fraction from its static properties Almirante and Urban (2024b, a, 2025); Almirante et al. (2025). Within these researches regarding the neutron effective mass or superfluid fraction, it was often the case that only the $^{1}\text{S}_{0}$ component was taken into considerations. However, in the context of astronomical simulations, the impact of spin-triplet superfluidity, such as the $^{3}\text{P}_{2}$ component, has been actively investigated Chau et al. (1992); Schwenk and Friman (2004); Page et al. (2004, 2009, 2011); Shternin et al. (2011); Haskell and Melatos (2015); Potekhin et al. (2015); Masuda and Nitta (2016, 2020); Leinson (2020); Masaki et al. (2020); Andersson (2021); Kobayashi and Nitta (2023); Marmorini et al. (2024); Granados et al. (2025); Nam and Sekizawa (2025); Hattori and Sekizawa (2025); Sedrakian and Rau (2025), and these studies are pushing their physical interpretation to the next stage. Although spin-triplet components are generally expected to emerge only in the neutron-star core region Sedrakian and Clark (2019); Wei et al. (2020), recent studies have begun to suggest their existence in the crustal matter or even in finite nuclei. For example, in Ref. Hinohara et al. (2024), it has been pointed out that, owing to the contribution of the spin-orbit interaction, $^{3}\text{P}_{1}$ superfluid components emerge in several open-shell nuclei. Another example is Ref. Tajima et al. (2023) where they revealed that spin-polarized neutron matter may produce the $^{3}\text{P}_{0}$ pairing component. There is a kind of neutron stars known as magnetars Makishima et al. (2014); Turolla et al. (2015); Kaspi and Beloborodov (2017), which possess surface magnetic-fields exceeding $10^{14}\,$G and possibly reaching up to $10^{17}$ or $10^{18}$ G Potekhin and Yakovlev (1996); Broderick et al. (2000); Bonanno et al. (2003); Naso et al. (2008); Frieben and Rezzolla (2012). In such matter with enormous magnetic-fields, the nuclear compositions and superfluid phase structures may be drastically converted Muzikar et al. (1980); Broderick et al. (2000); Peña Arteaga et al. (2011); Chamel et al. (2012); Basilico et al. (2015); Stein et al. (2016a, b); Yasui et al. (2019); Parmar et al. (2023); Sekizawa and Kaba (2023); Jiang and Chen (2024); Basilico et al. (2025), which should be thoroughly investigated based on microscopic theories.
@@
In our previous studies, under the title of “superfluid band theory,” we have pursued an approach in which superfluidity and band-structure effects are incorporated on the same footing into the density functional theory (DFT) Hohenberg and Kohn (1964); Kohn and Sham (1965); Nakatsukasa et al. (2016); Colò (2020) for the superfluid systems Yoshimura and Sekizawa (2024). There has been a wealth of literature where the DFT calculations successfully elucidated structures and properties of the inner crust matter, including pasta phases Magierski and Heenen (2002); Gögelein et al. (2008); Newton and Stone (2009); Pais and Stone (2012); Grill et al. (2014); Schuetrumpf and Nazarewicz (2015); Fattoyev et al. (2017); Schuetrumpf et al. (2019); Pęcak et al. (2024). For the superfluid extension of the nuclear DFT, we employ a local treatment of pairing interactions, known as superfluid local density approximation (SLDA) Bulgac and Yu (2002); Bulgac (2002); Jin et al. (2021). By extending the framework into systems at finite temperatures and under finite magnetic-field, we have further advanced the development of a tool to comprehensively describe the structures and properties of neutron star matter under extreme conditions, such as supernova matter and magnetars Yoshimura and Sekizawa (2025). In the present study, we extend our superfluid band theory to two-dimensional nuclear matter in the rod phase, which serves as a representative geometry of nuclear pasta. Our aim is to clarify how low-dimensional crystalline structures, in combination with spin-orbit interactions and strong magnetic-fields, influence neutron superfluidity, spin polarization, and pairing phases. To disentangle the roles of band-structure effects, magnetic-fields, and pairing correlations, we adopt a stepwise approach: we firstly analyze the entrainment effect within the Hartree-Fock (HF) framework (i.e., without pairing correlations), and subsequently incorporate superfluidity to investigate the interplay between spin-polarization and spin-singlet pairing, under the effect of the spin-orbit interaction. We then extend the superfluid density functional framework to include spin-triplet pairing interactions, and systematically examine the emergence of spin-triplet superfluid components by performing calculations with and without these interactions under various magnetic-field strengths. Through this unified and self-consistent framework, this work aims to provide microscopic insights into superfluid and magnetic phenomena in neutron-star crustal matter under extreme conditions.
@@
The article is organized as follows. In Section 2, the theoretical framework of this study is briefly explained, including the spin-triplet pairing. Section 3 indicates the computational settings used in the present work. The results of the calculations are shown and discussed in Section 4. Section 5 summarizes the work and suggests future prospects.
@@ 2 FORMALISM
In this section, we present the theoretical framework employed in this study, starting from the generic formulations of the Skyrme HF and Hartree-Fock-Bogoliubov (HFB) approaches. The formulation of the superfluid band theory has been developed in our previous work Yoshimura and Sekizawa (2024). For the spin-triplet pairing terms, detailed derivations can be found in Refs. Dobaczewski et al. (1984); Duguet (2014); Hinohara et al. (2024), although the notation used therein differs slightly from the present formulation. The extension to finite magnetic fields is based on the general formalisms discussed in Refs. Chamel et al. (2012); Stein et al. (2016a), and their implementation within our framework has been established in Ref. Yoshimura and Sekizawa (2025).
2.1 Skyrme-type energy density functional and Skyrme Kohn-Sham equations
The total energy per nucleon in the neutron star matter is, in general, written as
@@ 2.1 Skyrme-type energy density functional and Skyrme Kohn-Sham equations
where the nuclear part of the energy density is the sum of kinetic, interaction, Coulomb, and (for superfluid systems) pairing contributions,
$\mathcal{E}_{\text{nucl}}(\bm{r})=\mathcal{E}_{\text{kin}}(\bm{r})+\mathcal{E}_{\text{int}}(\bm{r})+\mathcal{E}_{\text{Coul}}(\bm{r})+\mathcal{E}_{\text{pair}}(\bm{r}).$ (2)
@@
$\mathcal{E}_{\text{kin}}(\bm{r})$ $=\sum_{q=n,p}\frac{\hbar^{2}}{2m_{q}}\tau_{q}(\bm{r}),$ (3)
@@
where the exchange term of the Coulomb energy is evaluated using the Slater approximation. For the interaction part, we adopt the Skyrme-type EDF which possesses the time-even and time-odd components, provided as
$\mathcal{E}_{\text{int}}(\bm{r})$ $=\mathcal{E}_{\text{even}}(\bm{r})+\mathcal{E}_{\text{odd}}(\bm{r}),$ (5)
$\mathcal{E}_{\text{even}}(\bm{r})$ $=\sum_{t=0,1}\big[C^{\rho}_{t}\rho_{t}^{2}+C^{\alpha}_{t}\rho_{t}^{2}\rho_{0}^{\alpha}+C^{\Delta\rho}_{t}\rho_{t}\Delta\rho_{t}+C^{\tau}_{t}\rho_{t}\tau_{t}+C^{\nabla J}\rho_{t}\gradient\cdot\bm{J}_{t}-C^{T}_{t}\sum_{\mu,\nu=x}^{z}(J_{t,\mu\nu})^{2}\big],$
$\mathcal{E}_{\text{odd}}(\bm{r})$ $=\sum_{t=0,1}\big[C^{s}_{t}\bm{s}_{t}^{2}+C^{s\alpha}_{t}\rho^{\alpha}_{0}\bm{s}_{t}^{2}+C^{T}_{t}\bm{s}_{t}\cdot\bm{T}_{t}+C^{\Delta s}_{t}\bm{s}_{t}\cdot\Delta\bm{s}_{t}-C^{\tau}_{t}\bm{j}_{t}^{2}+C^{\nabla J}\bm{s}_{t}\cdot(\gradient\times\bm{j}_{t}).$
The coupling constants $C_{t}^{\text{X}}$, where X stands for $\rho$, $\tau$, etc., are expressed in terms of the Skyrme force parameters, as summarized in Ref. Lesinski et al. (2007). Since the $\bm{s}_{t}\cdot\Delta\bm{s}_{t}$ term is known to induce spin instabilities Hellemans et al. (2012); Sekizawa and Yabana (2013), we set $C_{t}^{\Delta s}=0$ throughout this work. The local densities $\rho$, $\tau$, $\bm{j}$, $\bm{s}$, $\bm{T}$, and $J_{\mu\nu}$ denote the particle density, kinetic density, current density, spin density, spin-kinetic density, and spin-current tensor density, respectively. They are constructed from the single-particle wave functions as
@@
$J_{\mu\nu}(\bm{r})$ $=$ $\sum_{k,s,s^{\prime}}\frac{1}{2i}\quantity[\sigma_{\nu}]_{ss^{\prime}}\big[\psi^{*}_{ks^{\prime}}(\bm{r})\quantity(\nabla_{\mu}\psi_{ks}(\bm{r}))-\psi_{ks}(\bm{r})\quantity(\nabla_{\mu}\psi_{ks^{\prime}}(\bm{r}))\big],$ (11)
whrere $\bm{\sigma}=(\sigma_{x},\sigma_{y},\sigma_{z})$ denotes the Pauli matrices. The spin-current tensor can be decomposed into its scalar (trace), vector, and traceless symmetric components as
@@
where the mean-field potentials can be obtained from functional derivatives,
@@
$\frac{\delta E}{\delta J_{\mu\nu}}$ $=\mathsf{B}_{\mu\nu}(\bm{r}).$
Here we have omitted the isospin index $q$ for simplicity. These potentials are explicit functionals of the local densities, and the ground state is obtained self-consistently by iteratively solving the Kohn-Sham equations.